The Lowly Engineer
  • Metasys

    Gen Test, Challenges the OSC

    copy of email to Fred:

     

    Fred,

    Many problems at OSC today.  You guessed it Gen Test!  When they transferred back they took out a breaker for the 120V transformer that serves OSC.  Everything went down.  Now as the Metasys devices come back up, we are discovering everything moving way to slow.  All the OR reheat valves move very slow and one of the biggest problems of the morning was the steam valve for the heat exchanger.  It would not move but a percent every five minutes, so we lost all the OR’s and they had to delay Surgeries.   We got it all back up with overrides and monitoring.

     

    Do you think Kyle’s changes got blown out?   I also noticed that there was no reset schedule in, it was at 120* high and low.  I put the “output Low at 140” to try and help the temps come up with a little warmer water. (I actually had it at 160* but we overshot a couple rooms, so I bumped back down).

  • Study Material

    OSC Hybrid Room

    9/11/2014 10:49:04 AM  RM-P  Low Alarm 
    FE06B060 ATU53-16 – Hybrid OR #1 2580
    Value                            -0.0009 in wc
    Item Description                 Room Pressure

    Met with ACP yesterday morning to review options for a door closer on the door separating control room from the procedure room.  Expecting an installed closer by the end of next week.  Also notified dept manager, reminding/requesting that this door should be n.c..

  • Metasys

    INTERRUPTION OF REHEAT SYSTEM FOR OSC TUESDAY A.M

    Copied from an email notfication sent from Paul Dobry earlier:

    Hello: I wanted to make everyone aware that John Darr Mechanical (JDM) will be working on a punch list item in Mechanical Equipment Room #32 serving new Outpatient Surgery Center (OSC).  At or near 4 AM on Tuesday (8/26/14) JDM will secure power to the pumps serving the Terminal Unit (TU) system to perform some required re-piping.  During this time no hot water will be available for reheat throughout OSC.  SE on-duty should expect Metasys alarms for pump command not matching status.  Pumps will be re-started by 6 AM.  Bryan Diuble will be the point person for JDM and can be reached on his mobile phone if necessary at 734-216-5444.

     

    Please let me know if you have any questions or concerns.  Feel free to distribute to your teams as necessary.  THANKS!

  • Study Material

    Electrical Shutdown OSC & Surgery

    Copied from an email notification sent earlier: 

    To All,

      This is a reminder that at 4:30 am 8/26 all normal power (brown outlets) will be shut off for approximately 1 hour.  I will be on site and personally check with the Surgery and OSC staff prior to shut down to verify no case is under way.  I will leave my contact information in case an emergency case comes in so power can be restored right away.  The work to be done is very simple and I expect no issues.  Emergency power will be in operation the entire time.  Upon completion and restoration of normal power I will once again visit both Surgery and OSC to verify all systems are operating properly.

    Paul- Would you be so kind as to notify tonights powerhouse operator that the generators will run from roughly 4:30 to 5:30 am?  Thank you.

     Wm. E. Dreffs

    Ed Dreffs

    Project Engineer/Consultant

    Huron Valley Electric

    734-260-2746

    EDreffs@Huronvalleyelectric.com

     

  • Metasys

    OSC METASYS set points

    Worked with Fredasys and Kyle from JCI to improve METASYS response related to OSC chilled water and AHU 53-32’s d.a.t. set points.  It was noted by OSC staff that the procedure room’s humidity levels seemed high; although within our allowable range (20%-60%), it was determined that a better reset schedule that includes r.a. enthalpy would enhance the system’s performance.  Dr. Mazzeo should be happier…